Satellite Status

Report #5
Star Date: Nov. 2, 1998

The Comprehensive Performance Test (CPT) has been completed, and although there were the usual spate of minor software difficulties, it was a generally successful test. For the most part the hardware appears to be working properly, at least those systems that can be tested without being in vacuum.

Intense preparations are underway for the first pumpdown in the thermal-vacuum chamber. In the last 24 hours it was decided to do a very short (~half day) "pre-pumpdown" in an effort to release particulates that are thought to be adhering to the baffle walls. The idea is that as moisture on the walls evaporates, the particles will be released and fall to the mirror bench area. The mirrors will be covered, so after the chamber is brought back up to air, the bench will be carefully cleaned, the mirror covers removed, and the real thermal-vacuum pumpdown immediately begun. It is hoped that this will substantially decrease the particulate levels, which have been observed to increase when the instrument is bagged and purged, and therefore in a very dry environment.

There was some very good news this week. The collimator lamps on the LiF channels were briefly illuminated, and images were taken with the FES cameras on these channels. The alignment of the collimator to the instrument was found to be very good. This gives us hope that during the optical end-to-end test we will be able to locate the light sources quickly on the SiC channels, even though we don't have the benefit of FES cameras to assist us. (The FES cameras are only visible to the LiF channels.)

Reported by: Scott Friedman

(Click on photo to see enlarged version.)

Return to the List of Status Reports